OpenStudy (anonymous):

what is 0.58 in simpolest form

OpenStudy (anonymous):

29/50 or\[\frac{ 29 }{ 50 }\]

OpenStudy (texaschic101):

yep..because 0.58 = 58/100 and when reduced = 29/50

OpenStudy (anonymous):

Yes i knew it! lolz

OpenStudy (usman1995):

29/50